### v2

# Experimental Evidence for the Propagation and Preservation of Machine Discoveries in Human Populations

Abstract
Intelligent machines with superhuman capabilities have the potential to uncover problem-solving strategies beyond human discovery. Emerging evidence from competitive gameplay, such as Go, demonstrates that AI systems are evolving from mere tools to sources of cultural innovation adopted by humans. However, the conditions under which intelligent machines transition from tools to drivers of persistent cultural change remain unclear. We identify three key conditions for machines to fundamentally influence human problem-solving: the discovered strategies must be non-trivial, learnable, and offer a clear advantage. Using a cultural transmission experiment and an agent-based simulation, we demonstrate that when these conditions are met, machine-discovered strategies can be transmitted, understood, and preserved by human populations, leading to enduring cultural shifts. These findings provide a framework for understanding how machines can persistently expand human cognitive skills and underscore the need to consider their broader implications for human cognition and cultural evolution.
Machines powered by Artificial Intelligence (AI) have the potential to discover strategies that are difficult for humans to conceive: their computational capacity exceeds human processing power and speed; their ability to process and share information in serial or parallel with high fidelity allows them to distribute problem-solving efficiently [1]. AlphaGo, an AI system that defeated world champion Lee Sedol in the ancient game of Go, discovered superhuman strategies by playing millions of games against a copy of itself, and integrated these experiences into a complex strategy encoded in its neural network [2]. Given the human-defined objective of winning the game, the algorithm used extensive search to identify a successful strategy that has been described as ”alien”, considering thousands of years of human Go-play [3]. Recent evidence suggests that humans have significantly improved their own Go-play as a result of this innovation [4], while more subtle shifts that appear less ”alien” when viewed in historical context are also evident in Go opening moves [5]. In chess, selected strategic concepts extracted from an AI system were applied by grandmasters after observing the AI system only on a few examples [6]. The domain of gameplay remains, however, a rare example of human problem-solving strategies being shaped by machine-discovered strategies. This raises a question: Beyond being mere tools, under which conditions can machines persistently reshape the way we solve problems?
We address this question by integrating insights from the field of Cultural Evolution [7]: culture changes through a process of variation, transmission, and selection [8]. We propose that, correspondingly, machine-induced cultural shifts are contingent on the convergence of three critical conditions. Appropriate Discovery Difficulty: Strategies must be non-trivial and extend beyond the range of variations typically discovered by humans. Low Transmission Difficulty: Humans must be able to learn and effectively disseminate machine-introduced discoveries. Recognizable Selective Advantage: Cultural adoption of machine discoveries hinges on their demonstrated success and the extent to which humans attend to them and endorse them. As such, our proposal integrates strands of literature that have remained largely separate: prior research highlighting the transformative power of selective social learning [9, 10, 11]; observational evidence for the general possibility of machine-induced cultural shifts [2, 4, 6]; and conjectures about the complementary potential of human limitations and machine capacities [1].
In more detail, for a new problem-solving strategy to become part of the human cultural repertoire, it first needs to be discovered. Human discovery is limited by computational capacities and pervasive cognitive biases [12, 13, 1]. For example, human planning is generally confined to a few steps due to the combinatorial explosion of possible actions and outcomes [14]. In adapting to this challenge, humans use a range of heuristics [12], such as prioritizing immediate over long-term outcomes [15] or favoring familiar options [16]. In contrast, machines surpass humans in computational capacity, allowing them to rapidly accumulate experiences that would be too costly in terms of time, computation, or risk for humans to access. This may lead machines to discover strategies that are counterintuitive to humans yet have superior performance (see Figure 1B, vertical axis). Our experimental setup correspondingly reflects the cognitive heterogeneity of human-machine societies, thereby extending prior work which typically considers one agent type [11, 17, 18].
Figure 1: A. Transmission of machine discoveries can lead to cultural shifts. The multitude of potential transmission paths enables populations to preserve strategies that are difficult to transmit, fostering resilience against cultural loss. Machines, with their unique capabilities, may discover strategies that are improbable—or difficult to conceive—for humans. If such machine-discovered strategies are successfully transmitted and preserved within human populations, they could drive cultural shifts that outlast the presence of machines. B. Space of culturally preservable strategies. Human populations adopt strategies that are either easy to learn or sufficiently transmittable, forming a cultural frontier shaped by human limitations in time, computation, and communication (dashed line). We propose that some strategies, while practically inconceivable, may still be preservable by humans (shaded area). Machines, by discovering these strategies, could shift the cultural frontier (dotted line). Conversely, some machine-discovered strategies might not be preservable by humans, potentially fostering a reliance on machines (outer space).
Discovering superior strategies is not sufficient to induce a cultural shift, however [19]. The superior performance of the machine-discovered strategy must also be easy for humans to recognize as such [9, 20, 21, 22]. Furthermore, humans must be able to socially learn the strategy and transmit it to others [8]. The ability to select superior strategies and to transmit them in combination enables the population to maintain complex strategies that would be beyond individuals’ reach [11, 23]. When these mechanisms apply to machine-discovered strategies as well, machines could become sources of new knowledge that shift human culture (see Figure 1B, horizontal axis).
Even though intelligent machines are increasingly used in domains of human discovery, such as education and research [24, 25, 26], in most cases these AI systems do not convey problem-solving strategies that are entirely alien to human culture. As such, they do not satisfy the above conditions but remain within the confines of existing human discoveries (see Figure 1B bottom left). And even those AI systems that have discovered super-human knowledge, often remain – at present – tools in human hands rather than integrated into human cognitive frameworks. AlphaFold provides a remarkable example for this category: while it was inspired by the success of AlphaGo in emulating the intuition of human experts [27] and in doing so has achieved solving protein-folding problems at superhuman levels [28, 29], its discoveries have not (yet) transpired into novel human intuitions, unlike in the case of AlphaGo and AlphaZero. As such, it presents an unlearnable and unpreservable machine discovery (see Figure 1B top right).
Figure 2: A. Human-Machine Culture in the Lab. Participants were divided into populations of five Generations (0-4) with eight individuals each, including 15 Human-Machine Populations (three machines in Generation 0) and 15 Human-Only Populations. Lines represent a subset of demonstrator-learner connections (dashed: potential, solid: selected) involving a focal participant (dark magenta) in Generation 1 and its cultural ancestor (dark blue) and descendants (dark violet and grey). B. Experimental Stages. The ’Individual Learning Phase’ lasted five minutes, during which humans in Generation 0 managed to train on six networks, while machines managed to train on 1,000 networks. In the ’Social Learning Phase’, Generations 1-4 were informed about the performance of 5 randomly drawn players from the previous generation (demonstrator candidates), from which they selected their preferred demonstrator. They then observed and replicated the demonstrator’s strategy before solving the same network themselves and receiving feedback. In the ’Demonstration Phase,’ individuals recorded their strategy for the next generation. C. Network Task. Participants collected points on a network, with new potential connections appearing sequentially. We illustrate the two focal strategies in our task: maximizing immediate gains (myopic), and incurring initial costs for greater future rewards (optimal).
To provide the first experimental evidence of preserved, machine-induced cultural shift in humans, we designed a behavioral task that features conditions of appropriate discovery and transmission difficulty, along with a recognizable selective advantage for machine performance. The task emulated a prototypical situation in which human exploration is constrained by the pervasive human tendency to avoid losses (i.e., myopic decision-making). Human participants (n = 1,155) were paid to solve a series of “reward networks” (Figure 2 C), which we developed as a modified version of an established paradigm [30, 31]. In our task, networks consisted of nodes connected by links that were associated with rewards ranging from -50 to +400 points. Participants’ goal was to choose 10 moves through the network that would maximize the number of points gained. Our networks were designed such that there was an optimal strategy for solving them: incurring early losses subsequently led to links associated with the maximum reward of +400 points, which were not accessible when avoiding early losses. Hence, the optimal strategy for solving our networks contradicts the myopic tendency of humans to avoid losses and prioritize immediate gains, even when accepting short-term losses could be beneficial in the long run. In line with prior literature, we expected that this myopic tendency would make the discovery of optimal strategy in our task difficult for human players [32]. Meanwhile, we expected that a few minutes of training on our task would be sufficient for a deep-Q-learning algorithm (the ”machine player”) to successfully discover the optimal strategy. To compare the performance of human players on their own against the performance of human players exposed to machine players, we randomly allocated participants into populations that involved either only eight humans (15 ”Human-Only Populations”) or five humans and three machine players in their initial generation (15 ”Human-Machine Populations”). Each population consisted of five generations (numbered 0-4; Fig. 2 A), allowing us to observe differences between the two kinds of populations over time. After a series of individual attempts at solving networks, participants in Generation 0 solved four networks during the ”Demonstration Phase”, knowing that their moves would be recorded for future generations to learn from. Participants from Generations 1-4 could choose a ”demonstrator” from the previous generation to learn from, based on the demonstrators’ average performance. They did not know the demonstrators’ identity as a human or machine, but were informed that machine players were a possibility. Only after this Social Learning Phase did participants from Generations 1-4 demonstrate their moves for the next generation. All networks solved by the participants throughout the experiment were unique, requiring them to discover or learn a generalizable strategy rather than optimizing a specific solution.
In a second exploratory step, we sought to follow up on our experimental results with an agent-based simulation. Removing the feasibility constraints that apply to large-scale experiments with human participants, the simulation allowed us to systematically vary discovery and transmission difficulty, and to pitch random and selective social learning against one another [10]. In parallel to the experimental set-up, we organized agents into five generations, each consisting of eight agents. Agents could discover strategies through exploration or adopt strategies from the previous generation. Abstracting away from the task we used in the experiment, agents explored a trinary strategy space comprising the increasingly rewarding strategies: random, myopic, and optimal. For human-machine populations, three ”machine agents” with a discovery rate 1000 times higher than that of ”human agents” were placed in the first generation.
Humans Preserve the Machine Discovery
In our reward network task, achieving more than 2000 points was only possible by following the optimal strategy of incurring early losses to reach later links associated with the maximum reward of +400 points per move. Generally, human players did not find this optimal strategy on their own: Among the human populations, only one out of 600 participants discovered the optimal strategy over the course of the experiment, which four other human players then learned from them. In contrast, we found that the optimal strategy readily spread from machines to humans in the human-machine populations, resulting in the machine-discovery being preserved beyond the presence of the machines and across multiple generations: In 9 out of 15 Human-Machine Populations, human players learned this strategy in each of their generations. Correspondingly, we classified these populations as permanently preserving the machine discovery. In the remaining 6 populations, at least one human player in the generation immediately following the machine achieved more than 2000 points; but this discovery was only temporarily preserved as it did not transmit to later generations. Figure 3 illustrates prototypical cases for each of these qualitative categories. To summarise, the majority of Human-Machine Populations preserved the optimal strategy until the final generation in our experiment, while 14 out of 15 human populations did not discover the optimal strategy at all.
Figure 3: Four prototypical populations. We categorized populations based on the average scores of their top performers across generations. Scores above 2000 points indicate adoption of the optimal strategy. Among 15 Human-Machine Populations (panel A), 9 consistently exceeded this threshold, while the remaining 6 surpassed it temporarily in at least one generation following the introduction of machines. In contrast, only 1 of 15 Human-Only Populations (panel B) temporarily achieved scores indicative of the optimal strategy, while all the other Human-Only Populations did not. We illustrate one prototypical population for each case, with colors reflecting performance scores. Arrows represent demonstrator-learner relationships.
In terms of cultural lineages that emerged from demonstrator-learner relationships among the Human-Machine Populations (see Figure 3), we found that only machines had long-lasting impact on the cultural dynamics operating across generations: In the first socially learning generation of the Human-Machine Populations, most players (90%) selected a machine as their demonstrator, and by the last generation, every player was a descendant of a machine player. In contrast, none of the human players in Generation 0 of the Human-Machine Populations initiated a cultural lineage that persisted until the last generation.
Humans Benefit from the Machine Discovery
Figure 4: Evolution of Task Performance, Behavioural Congruency, and Strategy Description of Human Participants. Individual populations are represented as dots, with the grand average for each condition indicated by a diamond symbol. Human-Only Populations are indicated in green, Human-Machine Populations in orange. Performance (Panel A) refers to the average number of points earned during the Demonstration Phase. Bheavioural Congruency (Panel B) indicates the proportion of human player moves that a machine player would have made as well (contingent on making the same previous moves). Verbalised Strategy (Panel C) shows the fraction of human players who articulated a strategy referring to a deliberate acceptance of losses after the Social Learning Phase. Data points are jittered horizontally to improve visibility.
Panel A in Figure 4 shows the average reward gained in the task by generation and population. Descriptively, while the two types of populations started out with the same average in Generation 0, in Generations 1-4 Human-Machine Populations consistently earned higher rewards. Our preregistered linear mixed-effects model on these latter generations supported that this difference was significant ($\beta=-309.2$, SE = $57.0$, $\text{CI}_{95\%}$ = [420.7-420.7, 196.6-196.6]). Further models confirmed that Human-Machine Populations performed better than Human-Only Populations both in Generation 1, which could learn directly from machine players ($\beta=-404.6$, SE = $59.4$, $\text{CI}_{95\%}$ = [513.5-513.5, 293.0-293.0]), and in Generation 4, where influence from machines was remote by four iterations ($\beta=-229.1$, SE = $82.4$, $\text{CI}_{95\%}$ = [407.4-407.4, 72.3-72.3]).
Humans are Behaviourally Congruent with the Machines
Panel B in Figure 4 shows the proportion of human moves congruent with machine-generated moves, summarized by generation and population. Qualitatively, the results are similar to the results for rewards: no difference between conditions in Generation 0, but Human-Machine Populations are consistently more behaviourally congruent in Generations 1 to 4. A logistic mixed-effects model on these generations shows that Human-Machine Populations were more congruent than human populations ($\beta=-0.88$, SE = $0.14$, $\text{CI}_{95\%}$ = [1.14-1.14, 0.54-0.54]). Again, this was also the case within Generation 1, which learned from machine players ($\beta=-1.16$, SE = $0.14$, $\text{CI}_{95\%}$ = [1.43-1.43, 0.90-0.90]), and Generation 4 ($\beta=-0.66$, SE = $0.19$, $\text{CI}_{95\%}$ = [1.03-1.03, 0.24-0.24]).
Humans Internalize the Machine Discovery
During the Demonstration Phase of our experiment, players were asked to provide a written strategy of how they approached the task. In contrast to the demonstration itself, these written strategies were not transmitted to the next generation, and players were aware of this. Panel C in Figure 4 shows the proportion of human players mentioning the optimal strategy (i.e. the ”loss strategy”) in their written statements, as coded independently by three human annotators. In the Human-Only Populations, we observe practically no mention of the strategy, with proportions close to 0 throughout all generations; the exception to this is the single population that discovered the optimal strategy in Generation 3, resulting in 50% of players referring to it in Generation 4. In contrast, participants in the Human-Machine Populations consistently mentioned the optimal strategy explicitly, with proportions of about 30% to 50% per generation. A logistic mixed-effects model on Generations 1 to 4 shows that this difference is significant ($\beta=-4.18$, SE = $0.62$, $\text{CI}_{95\%}$ = [5.49-5.49, 3.04-3.04]).
Simulation Corroborates Conditions for Cultural Shift
Figure 5: Agent-Based Simulation Highlights Dependence on Discovery and Transmission Rates. An agent-based simulation replicates our five-generation social learning experiment, varying the difficulty of discovering and transmitting the optimal strategy. Panel A shows the 50% adoption boundary for the optimal strategy in the final generation, with lower discovery and transmission difficulty promoting its adoption, particularly under selective social learning (solid lines). Machines enable the adoption of hard-to-conceive but transmittable optimal strategies (orange line). Panel B illustrates the boost in average reward of mixed populations over human-only populations in the final generation, highlighting that this persistent boost is limited to strategies that are challenging for humans to discover but accessible to machines and that are of moderate transmission difficulty. Here we show the uplift for agents with selective social learning only.
We devised an agent-based simulation to systematically explore how variations in discovery and transmission difficulty shape the adoption of machine-discoveries and, ultimately, population-level success. Corroborating our experimental results, harder-to-discover strategies (with a higher (log(d)) were adopted when transmission difficulty 1t1-t was low, where $t$ is the transmission rate towards the end of the multi-generational simulation (Panel A in Fig. 5). This effect was enhanced when agents selected the highest-performing individual from the previous generation to socially learn from (solid lines) compared to learning from a random agent (dashed lines). When machine agents, with their increased discovery rate, were added, these human-machine populations could discover and preserve the optimal solution even when discovery was difficult (orange solid line). This machine-induced effect was not present when agents learned from a random agent of the previous generation (orange dashed line).
The ability of machine agents to discover hard-to-conceive strategies, combined with the ability of human agents to adopt and preserve these strategies, led to an increase in the rewards collected by human agents in the last generation of human-machine populations compared to their counterparts in human-only populations (Panel B in Fig. 5). We found that this uplift was constrained by both (a) high discovery difficulty, which renders human discoveries unlikely but machine discoveries likely, and (b) moderate transmission difficulty, which makes transmission feasible.
Discussion
Our work demonstrates that machine-discoveries can persistently reshape human problem-solving strategies when they are calibrated in difficulty, transmissibility, and selective advantage. In our experiment, superhuman computational capacity allowed machine players to discover a superior strategy that offset human cognitive biases toward myopic and suboptimal solutions and, as such, was effectively beyond human reach. This does not mean that the optimal strategy was impossible to discover for human participants at all, but that fundamental limitations in human cognitive resources, attention, and time [1] made it exceedingly rare. Yet, the machine-discovered strategies were still learnable for human players, who did not merely replicate them but also applied them to new variations of the task. Further signifying the integration of the counterintuitive problem-solving strategy into their cognitive framework, many human players explicitly articulated the strategy in writing. The machine-discovery’s selective advantage translated into human players overwhelmingly favoring machine over human ancestors, producing persistent cultural lineages where every human player in the last generation was ultimately a descendant of a machine player in the first generation.
It is interesting to consider these results in view of prior work from cultural evolution, which has shown that large effective population size can effectively buffer against risks of losing complex discoveries [33, 17]. This suggests that the scope of machine-induced cultural shifts may expand in contexts where large groups interact with machine-generated strategies. For instance, the widespread adoption of generative AI in educational, professional, and creative contexts naturally exposes millions of people to machine-mediated strategies for creation or problem-solving. Larger population sizes increase the probability that machine-discovered strategies are learned in the first place, as a diverse range of individuals with varying levels of skills engage with them. Analogously, skilled individuals can prevent strategies from being lost, allowing less skilled individuals to benefit from more demonstrations and eventually acquire the strategies [17], thereby ultimately accumulating a critical mass of individuals for triggering a tipping point in their cultural adoption [34]. Emergent evidence suggests that even simple machine agents could further boost these group dynamics’ efficacy [35].
Yet another factor that could broaden the scope of machine-induced cultural shifts is pedagogical behavior. Teaching is one of the primary mechanisms through which humans transmit knowledge and skills [8, 36], and experiments have shown that the value of teaching increases with the complexity of strategies [37]. Notably, discovery and transmission difficulty have also been identified as key determinants in the evolution of teaching [38], underscoring that machine-induced cultural shifts may occur within a relatively narrow range of conditions. This line of work does, however, also point towards another possible pathway that could enhance the efficacy of machine-induced cultural shift: more explainable AI systems could provide both teachers and learners with insights that facilitate the adoption of machine-discovered strategies, and enhance their scope through more effective human-machine interactions [39]. A deeper understanding of social learning between humans and machines across a wider range of experimental paradigms and applications is decisive to identifying the conditions under which machine-discovered strategies are most likely to be adopted and transmitted within human groups, thereby shaping cultural trajectories in meaningful ways.
A major question arising from these results concerns the scope of machine-induced cultural shifts. Our pilot experiments illustrated the key challenge of achieving the right balance between discovery difficulty and transmissibility: At one extreme, simple strategies were readily discovered by humans without machine assistance, limiting their capacity to induce meaningful shifts. At the other extreme, strategies requiring lengthy and precise sequences of actions were too complex for human participants to learn and pass on. It is remarkable that the optimal strategy in our final design was discovered by none of the 195 human participants in the first generation (in deviation from previous work such as [11], where discovery-rate in the first generation was still at 13% ). Nonetheless, in our setting, this optimal strategy ultimately still was learnable and preservable as soon as it was discovered.
The work presented here suggests that by learning from machines, humans may adopt and transmit hard-to-conceive optimal strategies, thereby broadening humans’ cognitive repertoire without incurring the costs typically associated with discovering such strategies. Our results apply particularly to scenarios in which human exploration is constrained by time rather than computational cost, machine exploration remains scalable. For instance, chemical recipes can be explored at high throughput and scale through automation [40]. Future research could examine whether the extensive world knowledge embedded in large generative models might similarly facilitate discovery processes, even in contexts where exploration is costly and the data available to both humans and machines is comparable.
Theoretical frameworks about cumulative cultural evolution posit that culture evolves through the iterative refinement of adaptive solutions [41], implying that certain outcomes may remain undiscoverable without intermediate adaptive steps. Future studies could explore how the enhanced exploratory depth and breadth enabled by machines might help overcome these limitations imposed by structural factors, such as entrenched path dependencies and biased expectations shaped by a historically conditioned epistemic horizon [42, 26]. Supporting this conjecture, recent work has highlighted machines’ potential to identify novel strategies for fostering cooperation that humans have proven unlikely to consider [43]. Even when it comes to changing notoriously entrenched conspiracy beliefs, intelligent machines (here, LLMs) are emerging as potent agents of persistent learning among humans [44]. But these contexts also point to the potential challenge that machine-discovered strategies might sometimes be undesirable from a human perspective and not be explored for good reasons, such as ethical concerns or social acceptability. Relatedly, human adoption and preservation of machine-discovered strategies may raise ethical concerns about human autonomy and creativity as much as it might bear epistemic risks that could undermine human knowledge [45]. In other contexts, AI aversion or other (oftentimes variable) perceptions of machine-derived knowledge [46, 47, 31] might influence adoption rates. However, as it becomes increasingly difficult to dissociate human- from machine-derived knowledge, it is all the more crucial to understand how machines might expand the frontiers of human thought and enable the emergence of culturally significant, hard-to-conceive strategies.
Methods
Participant Recruitment, Exclusion and Compensation
Participants (N = 1,155) were recruited through the online platform Prolific [48], while screening for individuals residing in the United States and the United Kingdom, being proficient in English, and maintaining an approval rating of at least 95%. Upon successful completion of the study, participants received a fixed payment of £2.25, complemented by performance-dependent bonus compensation payments averaging £0.70. The median time for completing the study was 13:45 minutes. The average hourly compensation rate was £9.82.
As preregistered, participants were excluded based on two criteria: 1) Participants who took more than 25 minutes, deemed unreasonably long; 2) Participants with missing moves in any of their demonstration trials, as subsequent generations could not replicate incomplete solutions. These criteria also implicitly served as an attention check. Of the 1,303 participants who completed the experiment, 148 were excluded under these conditions. Their spots were then filled by newly recruited participants. No other exclusions were made.
Participants who passed the exclusion criteria were eligible for bonus compensation. Participants’ compensation depended in part on their own performance in the Demonstration Phase, and in part on the performance of people that learned from them. To ensure a fair chance of gaining a bonus payment independently of the condition a participant was sampled in, we ranked these scores separately per generation and condition. The final bonus was then computed as the percentile score rank multiplied by £1.40.
Network Task
We adapted the Reward Network task [30, 31] consisting of a network $g$ with 12 nodes and 30 edges, where each edge is associated with a reward $r\in[-50,0,100,200,400]$ (see figure 2 panel C). Hidden to participants, each node was associated with a level $l\in[0,1,2,3]$. Constraints in the network generation ensured that any path connecting a node of level 0 to a node of level 3 included at least three edges with negative rewards. The largest reward of 400 points was accessible only from level 3. These two constraints required participants to accept at least three losses to access the highest rewards. Transitions between nodes of level 0 contained the second-highest reward of 200 points, making myopic behavior a reasonable but generally suboptimal strategy.
Participants began the game at a predetermined starting node (e.g. node B in figure 2 panel C). Initially, only the outgoing edges from this node were revealed. As participants traversed the network, they sequentially uncovered it and the outgoing edges of each visited node were displayed. Each participant would be playing on, in total, 10 different networks drawn randomly from the pool of experimental networks. The assignment of networks to individual participant trials ensured that no participant experienced the same network twice.
We generated three sets of networks (training, validation, and experiment) with 1,000 networks each. The first two sets were used only during the training of the neural policy, while the last set was used exclusively during the experiment. After generation, we computed scores for a myopic rule-based strategy and a loss-seeking rule-based strategy. The myopic strategy always selects the move with the immediate highest reward, while the loss-seeking strategy prioritizes edges with a -50 reward over all others. We excluded networks from all three sets where the myopic strategy outperformed the loss-seeking strategy, which accounted for approximately 9% of all generated networks.
Populations and Social Learning Opportunities
The experiment consisted of 30 populations, each with 5 generations and 8 players per generation (see figure 2.A). Half of the populations were assigned to the ”Human-Machine” condition, and the other half to the ”Human-Only” condition. In the ”Human-Machine” condition, machines took three player positions in the first generation. Starting from the generation 1, each participant could choose one of five possible demonstrators drawn from the eight positions in the previous generation. Participants were dynamically assigned to available positions in the populations, prioritizing the completion of existing populations before adding participants to new ones. Positions for later generations became available only after all potential demonstrators had completed the experiment.
Preregistration
The study was preregistered on asPredicted.org, including a detailed description of predictions, measures, conditions, analyses, exclusion criteria and sample size (https://aspredicted.org/3C3_HJH). In particular, we preregistered the following predictions:
  1. 1.
    Main predictions
    1. (a)
      General benefit of AI. Overall, participants from generation 1 onwards in AI trees will outperform participants in human trees.
    2. (b)
      Long-lasting benefit of AI. The subset of participants in the final generation of AI trees will outperform the final generation of the human trees.
  2. 2.
    Secondary predictions
    1. (a)
      Transmission from AI to humans. The subset of participants in generation 1 of the AI trees will outperform generation 1 in the human trees.
    2. (b)
      Explicit recognition of counterintuitive strategy. Overall, participants from generation 1 onwards in the AI trees will describe the counterintuitive strategy more frequently in their written strategies than those in the human trees.
Experimental Flow
Panel B in Figure 2 depicts the main phases of the experiment. A static version mimicking the experience for participants in later generations is available here:
Below, we summarize the main phases of our experiment:
  1. 1.
    Introduction Phase: Participants learned about the study’s purpose, procedures, data handling, estimated time, and financial compensation. They then completed a tutorial guiding them through the task and interface.
  2. 2.
    Individual Learning Phase: Participants solved the task individually. Generation 0 worked through 6 networks, while later generations worked through only 2 networks (given they encountered four more in the Social Learning Phase, see below) to familiarize themselves with the task. See also Figure S1.
  3. 3.
    Social Learning Phase (only from generation 1: Participants selected a demonstrator (see Figure S2). They could base their decision on demonstrators’ average scores from the Demonstration Phase (they were also informed about their own average score as a reference point). They then went through four networks by observing a replay of the demonstrator’s solution (see Figure S3), repeating that solution (see Figure S4), and trying to solve the same network on their own. Last, they received feedback comparing their solution with the demonstration.
  4. 4.
    Demonstration Phase: Participants independently solved four networks. The solutions entered were used for the Social Learning Phase of the next generation for those selected as a demonstrator.
Participants also documented their personal strategy in written form, before and after the Social Learning Phase.
Neural Policy and its Training
We used a deep Q-learning-based reinforcement learning approach to train a neural policy for navigating and solving reward network tasks, as outlined by [49]. The neural policy featured a gated recurrent unit (GRU) [50], flanked by two linear layers with ReLU activation, each containing 15 hidden units. The input $Q(o_{t})$ consisted of a one-hot-encoded reward of each target node in a vector of shape $Nodes\times Rewards$}, while the output represented the Q-value of each target node, with unreachable nodes set to negative infinity. Thus, similar to the initial experience of human participants, the algorithm could only process the immediate outgoing edges from the current node without engaging in any explicit planning.
The neural network approximated Q-values $Q(o,a;\theta)$ where $\theta$ represented the network parameters. These parameters were optimized by minimizing the loss function:
$L(\theta)=\mathbb{E}\left[\left(y_{t}-Q(o,a;\theta)\right)^{2}\right]\text{,}$
with the target value $y_{t}$ defined by the Bellman equation:
$y_{t}=r+\gamma\max_{a^{\prime}}Q(o^{\prime},a^{\prime};\theta^{-})\text{,}$
and $\gamma=0.99$ as the discount rate. The target policy weights $\theta^{-}$ were only updated every 200 steps to stabilize training.
An epsilon-greedy strategy guided training with $\epsilon$ starting at 1 and reducing according to $\epsilon_{e}=0.99^{\lfloor e/1,000\rfloor}$, reaching a minimum of 0.01. Trajectories were stored in a replay buffer of 500 episodes, and a batch of 16 episodes was sampled for updating the policy. The policy was optimized using the Adam optimizer with an initial learning rate of $1\times 10^{-3}$, adjusted downwards by a factor of 0.8 every 2,000 episodes. The agent underwent 20,000 episodes of training, with evaluations every 100th episode using 1,000 test networks from a hold-out set. Training took approximately 5 minutes on an RTX 5,000 GPU. We depict the average performance on a hold-out test set in Figure S5, illustrating that the policies initially discovered the myopic strategy before eventually settling on the optimal strategy.
Measures
Task Performance
We measured task performance by calculating the total score for each trial, which involved summing the rewards from the 10 moves. The maximum achievable reward was 2,650, which corresponded to taking three losses of 50 points each and then gaining 400 points seven times. In contrast, the maximum score attainable with a myopic strategy was 2,000, yielded by ten gains of 200 points each. The exact scores achievable with these two strategies varied between networks.
Behavioural Congruency
We assessed human-machine behavioural congruency by comparing the actions taken by humans, denoted as $a^{H}_{l,t^{\prime}}$, with those chosen by the machine
$a^{M}_{l,t^{\prime}}=\arg\max_{a}Q(m^{H}_{l0},...,m^{H}_{lt^{\prime}})\text{,}$
based on the trajectory of moves $m^{H}_{lt}$ selected by the human participant. A match $a^{H}_{l,t^{\prime}}=a^{M}_{l,t^{\prime}}$, where the machine’s action corresponded with the human’s action, was encoded as ’1’; a mismatch as ’0’. This method required processing the entire trajectory to ensure that the recurrent neural unit of the neural policy could utilize the accumulated historical information to influence its decisions.
Written Strategies
To measure the explicit recognition of the loss strategy, we manually coded all 2,310 strategy descriptions (covering both time steps for every participant) with the aid of three trained human annotators. The annotators familiarized themselves with the task and independently coded the written strategies using a binary variable to indicate the presence of the loss strategy. Before coding the full experimental data, the annotators calibrated their ratings by coding a small independent dataset (n = 76) from two previous pilot studies. After individual coding, interrater agreement was high (Light’s $\kappa=0.83$). They then discussed any discrepancies before proceeding to code the full dataset. Annotators were blind to participant conditions during both calibration and full data coding, and data was randomly shuffled prior to coding to prevent any bias linked to participant order. For the full dataset, interrater agreement was excellent (Light’s $\kappa=0.92$).
Statistical procedure
All analyses were performed using the lme4 package (version 1.1-35.1; [51]) in R (version 4.3.2; R Core Team), and the structure and procedure of these analyses were part of our preregistration. We first subset the data to only include demonstration trials of participants. Next, we aggregated the data to the appropriate level for each measure of interest: trial-level for reward, move-level for behavioural congruency, and participant-level for the written strategies.
In all models, the predictor of interest was a fixed effect of condition. For the models predicting the measure across generations 1-4, we included random intercepts for participant and population, along with a random slope for generation within population. For models predicting measures within individual generations, we incorporated the random intercepts for participant and population. Reward outcomes were modeled using a linear link function, whereas behavioural congruency and strategies were modeled using a logistic link function. As preregistered, we centered and scaled the variable generation to enhance model convergence and to make the main effect of condition more interpretable.
Power analysis
We determined the goal of 15 trees per condition by simulating data via an agent-based model resembling a variant of the model described in the previous section. In this model, rates of learning success and individual discovery rates where calibrated based on pilot data. In a power analysis on this simulated data, we resampled runs with 15 trees each 1000 times and fit the model for prediction 1b (since it has the smaller subsample of the two main predictions). We found our predicted effect, as assessed by a t-value greater than 2 for ”condition”, in 957 ( 96%) of the samples, suggesting sufficient power to assess all predicted effects.
Agent-based simulation
We simulated a scenario analogous to our experiment using an agent-based simulation. As in our experiment, agents were organized into $G=5$ generations, labeled $g=0,1,2,3,4$. Each generation consisted of $N_{\text{gen}}=8$ agents. To constrain the simulation, we selected most parameters to loosely resemble those of our experiment, focusing exclusively on exploring the consequences of transmission difficulty and discovery difficulty on the behavior of agents in the last generation ($g=4$).
In the simulation, each task $i$ could be solved using one of three discrete strategies:
$\mathcal{S}_{i}=\{S_{\text{random}},S_{\text{myopic}},S_{\text{optimal}}\}.$
The reward for completing a task with a given strategy was computed as
$R_{S}\sim\bar{R}_{S}+\mathcal{N}(0,\sigma^{2}),$
where $\bar{R}_{S}$ was the average reward for the strategy and $\mathcal{N}(0,\sigma^{2})$ represented Gaussian noise with variance $\sigma^{2}=200$. The average rewards were $\bar{R}_{\text{random}}=600$, $\bar{R}_{\text{myopic}}=1400$, and $\bar{R}_{\text{optimal}}=2200$.
Each agent had a single trinary state, representing its preferred strategy, which was initialized with the random strategy ($S_{\text{random}}$). In the initial phase, Individual Exploration, agents either used their preferred strategy or, with equal probability $\epsilon_{\text{exp}}=0.5$, explored a new strategy. During exploration, the discovery rates $d_{\text{optimal}}\in[10^{0},10^{-6}]$ and $d_{\text{myopic}}=0.4$ determined the likelihood of discovering the optimal and myopic strategies, respectively. Upon discovering a new strategy, the agent updated its preferred strategy if the new strategy was superior to the previous one. To mimic a fundamental shift in processing speed for machines, the discovery rate for machine agents was set to be 1,000 times higher than that for human agents.
In parallel with our experiment, agents in the first generation ($g=0$) solved six tasks through individual exploration. This phase was followed by a Demonstration phase, in which agents completed four tasks exclusively using their preferred strategy, without exploration. For agents in subsequent generations ($g=1,2,3,4$), each agent completed two tasks through individual exploration, followed by four tasks during a Social Learning phase.
During social learning, each agent was randomly assigned five demonstrators from the previous generation. We used two modes of demonstrator selection: either, (a) randomly selecting an agent (Random Social Learning), or (b) selecting the agent with the highest average reward in the demonstration phase (Selective Social Learning). In the subsequent four social learning tasks, agents imitated the strategy of their demonstrator with a probability determined by the transmission rate ($t\in[0.70,1]$) and updated their preference state correspondingly.
We conducted the simulation for two population configurations: one consisting solely of human agents and another including three machine agents in the first generation. The simulations were run across a grid of discovery difficulties ($\log{d_{\text{optimal}}}$) and transmission difficulties (1t1-t) with 100 replications for each parameter combination. As a robustness check, Supplementary Figure S11 includes results for simulations with a single machine in generation 0, participant selection from all previous-generation agents, and an increased population size of 16.
Acknowledgments
We thank Jaeeun Shin and Rodrigo Schettino for their support in data collection. We also thank Peter Dayan and Wolfgang Hönig for their valuable feedback.
Author contributions:
L.B., T.F.E., M.D., and I.R. conceived the research project and experimental methodology; L.B., T.F.E., A.-M.N., S.B., and V.C. refined the methodology; L.B., T.F.E., A.-M.N., M.D., and I.R. developed the theoretical grounding; T.F.E. pre-registered and conducted the formal analysis; L.B. conducted the agent-based simulation; S.B. developed the machine player; V.C. implemented the experiment; L.B., T.F.E., and A.-M.N. administered and supervised the study; L.B.,T.F.E., and A.-M.N. created the visualizations; L.B., T.F.E., A.-M.N., M.D., and I.R. wrote the paper.
Competing interests:
There are no competing interests to declare.
Data and materials availability:
All data and code for reproducing the statistical analyses, visualizations, training of the machine player, and execution of the experiment are available on OSF [52], including a static version of the experimental interface [53]. A preregistration document is available on AsPredicted [54].
Supplementary information
Supplementary Text
Figs. S1 to S10
Tables S1 to S3
Supplementary Information for
Experimental Evidence for the Propagation and Preservation of Machine Discoveries in Human Populations
Levin Brinkmann, Thomas F. Eisenmann∗†, Anne-Marie Nussberger,
Maxime Derex, Sara Bonati, Valerii Chirkov, Iyad Rahwan
Corresponding authors. Email: eisenmann@mpib-berlin.mpg.de (T.F.E.); rahwan@mpib-berlin.mpg.de (I.R.)
These authors contributed equally to this work.
This PDF file includes:
Supplementary Text
Figures S1 to S10
Tables S1 to S3
Supplementary Text
Classification of populations
Figure S9.D displays the average points of the highest performing participant in each population and generation. In the reward network task, a point value above 2000 is only possible if the largest level 3 is reached and multiple largest rewards of 400 points are collected. Therefore, we classify a generation of a population as ’discovered’ if the average point value of the top performer exceeds 2000 points. Focusing on the last generation of each population, we found 9 groups exceeding this threshold in the human-machine population and one group in the human population.
Correspondingly, we found:
  • Permanently preserved: In 9 out of 15 human-machine populations, humans permanently preserve the machine discovery.
  • Temporarily preserved: In 6 out of 15 human-machine populations, humans temporarily preserve the machine discovery. This means that the machine discovery was preserved at least in the second generation; however, no participant in the last generation exceeded the threshold.
  • Discovered: In 1 out of 15 human populations, the optimal strategy was discovered, and the maximum average point value in the last generation exceeds 2,000 points.
  • Not discovered: In 14 out of 15 human populations, no participant exceeded an average point value of 2,000.
Qualitative summary of prototypical strategies
Many written strategies referred to “looping” as a feature of the task that was related to higher performance, but not necessarily optimal: ‘the trick is to identify a path between two nodes that both award high points and then repeat the path back and forth.’, ‘Find two letters with connecting green arrows and go back and forth between them.’ Some participants admitted that they did not have a good plan of what to do: ‘no idea seriously just choosing patterns’, ‘Unsure of an effective strategy’. Others confidently reported “superstitions” that focused on parts of the task that had no actual relation to good performance: ‘remember the first random letters then remember a sequence’, ‘Longer routes seemed to provide more options for success’.
Frequently, written strategies explicitly referred to myopic behavior: ‘Ignoring the big loss moves’, ‘Always avoiding neutral or negative paths seems to be the best strategy.’ Sometimes, these myopic strategies even explicitly highlighted participants’ bias against early losses: ‘always choose the line that gives you points instead of taking away points for less risk’, ‘I just followed the green paths, I don’t even really know what else I would try to do? Picking the negative point options makes no sense, after all.’ In contrast, participants that had learned the loss strategy frequently referred to it explicitly and showed their long-term planning: ‘Take a loss for the first three moves then move between two high scoring nodes.’ ‘if you take a hit early (eg the first 3 nodes are -50), you are in a position to maximize the following 7 nodes, each with +400. i tried it differently initially, but player 2’s method is best.’ The last example also emphasizes the important role of social learning for developing this participant’s strategy.
Last, the one exceptional human player in the human populations that discovered the loss strategy on their own wrote ‘I try o track my best point’ (sic) after training, maybe indicating an approach that was open to exploring while systematically assessing their point score. After social learning, they wrote ‘I tried to keep losing till I see the thinker green line.’ (sic), clearly indicating the loss strategy even though no previous participant was able to demonstrate it to them. Presumably, the participant made good use of the “try yourself” trials to explore also unusual options, and discovered the loss strategy this way; exceptionally so, since they are the only player to do so in the 15 human populations.
Task performance
Figure S9.A shows the average reward per demonstration trial, summarized by generation and population. Descriptively, human participants in the average human population started out equal to the average human-machine population in generation 0, but human-machine populations earned consistently higher rewards in generations 1 to 4.
We ran the linear mixed-effects model testing prediction 1a (”General Benefit of AI”) on the demonstration trials of generations 1 to 4, with total reward as the dependent variable and fixed effects for condition, generation, and their interaction. Additionally, we included random intercepts for participant and population, as well as a random slope for generation on population. As preregistered, we centered and scaled generation to improve model convergence and make the main effect for condition interpretable. In line with our prediction, the model revealed a significant main effect for condition ($\beta=-309.2$, SE = $57.0$, $\text{CI}_{95\%}$ = [418.4-418.4, 200.1-200.1]), indicating that human-machine populations outperformed human populations overall.
The model for prediction 1b (”Long-lasting Benefit of AI”) followed a similar logic, but was using only the data of generation 4. Given the more restricted data set, only the fixed effect for condition and the random intercepts for participant and population remained. Similar to the first model, there was a meaningful main effect for condition ($\beta=-229.1$, SE = $82.4$, $\text{CI}_{95\%}$ = [383.1-383.1, 55.5-55.5]), supporting our prediction that human-machine populations would perform better than human populations by the last generation.
The model for prediction 2a (”Transmission from AI to Humans”) was identical to the one for 1b, but used the data of generation 1. Again, there was a significant main effect for condition ($\beta=-404.6$, SE = $59.4$, $\text{CI}_{95\%}$ = [524.8-524.8, 286.1-286.1]), supporting our prediction that human-machine populations performed better than human populations in the immediate generation learning from machine agents. Here, the model returned a warning about a singular fit, so we also checked all other optimizers available in lme4; however, since all optimizers converged to identical values for the fixed effects, we concluded the warning to be false-positive.
Human-Machine Behavioural Congruency
Figure S9.B shows the proportion of human moves congruent with machine-generated moves per demonstration trial, summarized by generation and population. Qualitatively, the results are similar to the results for rewards: no difference between conditions in generation 0, but there is consistently more behavioural congruency in human-machine populations for generations 1 to 4.
The models assessing predictions 1a to 2a were identical to the ones for reward, with exception of the different outcome measure and being on the level of moves per trial. Since behavioural congruency was a binary variable, we used a logit link here. All three models replicated the expected effects for condition found for reward on the behavioural congruency measure: Human-machine populations were behaviourally more congruent than human populations overall ($\beta=-0.88$, SE = $0.14$, $\text{CI}_{95\%}$ = [1.16-1.16, 0.59-0.59]), in the last generation ($\beta=-0.66$, SE = $0.19$, $\text{CI}_{95\%}$ = [1.02-1.02, 0.31-0.31]), and in the first generation ($\beta=-1.16$, SE = $0.14$, $\text{CI}_{95\%}$ = [1.42-1.42, 0.90-0.90]). Again, the last model returned a warning about a singular fit, but all available optimizers converged to identical fixed effects until the second digit, so we took it as a false positive.
Written Strategies
Figure S9.C shows the proportion of human participants reporting the loss strategy, summarized by generation and population. In the human populations, we observe practically no mention of the loss strategy, with proportions close to 0 throughout all 5 generations; the exception to this is the single population that discovered the loss strategy towards the end, resulting in 50% of the participants referring to the strategy in generation 4. In contrast, participants in the human-machine populations consistently mentioned the loss strategy explicitly, with proportions of about 30% to 50% per generation. In generations 1 to 4 (i.e. with social learning), almost no participants mentioned the loss strategy immediately after the training trials (human populations: 1.7%; human-machine populations: 2.7%), but 39.8% of participants in the human-machine populations mentioned it after social learning had occurred (compared to 1.9% in the human populations). In comparison, in generation 0 (without social learning) very few participants in either condition referred to the loss strategy at either time point (human-machine t1: 1.3%; t2: 1.3%; human-only t1: 2.5%; t2: 5.8%).
The model assessing prediction 2b (”Explicit Recognition of Counterintuitive Strategy”) was identical to the one for prediction 1a, but using a logit link due to the binary outcome variable and missing the random intercept for participants, since it was built on participant level. Again, there was a significant main effect for condition ($\beta=-4.18$, SE = $0.62$, $\text{CI}_{95\%}$ = [6.30-6.30, 3.11-3.11]), indicating that human-machine populations explicitly mentioned the loss strategy more frequently than human populations from generations 1 to 4. This model also returned a warning about a singular fit, but all available optimizers converged to qualitatively similar parameters.
Agent-Based Simulation
We run an agent-based simulation as described in Supplementary Materials and Methods for a grid of learning rates $l$ and discovery rates $d$. Specifically, we use 36 linearly spaced learning rates ranging from 0 to 0.35, and 25 logarithmically spaced discovery rates ranging from $10^{-6}$ to 1.
In Supplementary Fig. 5, panels A and B, we illustrate the average reward of human and human-machine populations, respectively. These simulations reflect the experimental condition in which players could choose one of five possible demonstrators from whom to learn.
For human populations, when the discovery rate is high, the optimal solution is rediscovered in each generation, resulting in an average reward that remains largely independent of the social learning rate. In contrast, when the discovery rate is very low, no human population discovers the optimal solution. We observe an intermediate range where social learning enhances the average reward. Within this range, the average reward interacts with both the social learning rate and the discovery rate, with a high social learning rate enabling populations to retain rarely discovered solutions.
In human-machine populations, machines consistently identify the optimal strategy. Under sufficiently high social learning rates, this solution is preserved across the five simulated generations. Accordingly, we observe that for discovery rates in the range of $10^{-3}$ to $10^{-5}$, average rewards increase at low to medium social learning rates.
In Supplementary Fig. 5, panels C and D, we present the same metrics for populations where demonstrators were selected randomly. Here, the adoption of optimal solutions occurs only at lower discovery rates compared to populations that selected their demonstrators (compare Panel A to C). More strikingly, the maintenance of solutions discovered by machine players is almost completely absent when populations engage in random social learning (compare Panel B to D).
Pilots
We run 10 cycles of pilots. Each in each cycle we run some of a set four different types of pilots indicated by the letters A to D. These type of pilots differ in the experimental setup and therefore allowed to test different premises.
  • Pilot type A: Participants are placed in generation 0 of the experiment. They therefore have no chance of social learning. This pilot allows conclusion about the discoverablity of the loss strategy.
  • Pilot type B: Participants are placed in the first generation of the experiment. Solutions of the first generation are provided by rule based algorithm. Generally, we provided participants with a mix of teachers. Some of the teacher would show a solution of the loss strategy while other teacher would display a myopic solution. This pilots allows to infer about the learnability of the solution, provided the participant is exposed to the loss strategy. This type of pilots also allows to infer about the selection of a teacher by the participant.
  • Pilot type C: This pilot type was similar to pilot type B, however, participant could only choose from myopic solutions to learn from.
  • Pilot type D: This pilot type was the most complex of all. Here we collected data from the 0th and 1st generation. No machine solution was provided. Like pilot A, this pilot was intended to investigate the learnability. Additionally this pilot allowed a technical check of the multi-generational assignment.
Not all pilots where informative. In the following we present a selection of all pilots.
  • Pilot 1 A / B
    • Setup: The network task presented in these earlier pilots is similar to the final reward network task, with important differences:
        The full network is shown to participants, but humans are expected to behave suboptimally due to cognitive constraints in exploring the full solution space.
        The network is smaller, with only 8 nodes.
    • Results: Participants fail to discover the loss strategy. Even when shown the loss strategy, they fail to repeat it.
  • Pilot 2 A / B
    • Setup:
        Increased the time for participants.
        Made it easier for participants to reproduce the correct strategy by showing it in written form.
        Asked participants to write down their strategies.
    • Results:
        In Pilot 2 A, participants enacted and described a myopic strategy.
        In Pilot 2 B, only 4 out of 10 participants were able to reproduce the loss strategy, either in practice or verbally, even after explicit instruction.
  • Pilot 3 B
    • Setup: Focused on the learnability of the loss strategy:
        Intensified the social learning phase.
        Introduced iterative observation and try-yourself pages, allowing participants to play the task themselves and compare their performance to the teacher’s.
    • Results: Only 2 out of 10 participants somewhat learned the loss strategy.
  • Pilot 4 B
    • Setup: Made the written correct strategy even more salient.
    • Results: No significant improvement in learnability.
  • Pilot 5 B
    • Setup:
        Simplified the task.
        Redesigned the networks to better separate the myopic and loss strategies in terms of point values.
    • Results: Most participants learned the loss strategy and were able to reproduce it in demonstration trials.
  • Pilot 5 A / C
    • Setup:
        Conducted two more pilots with the same experimental design as Pilot 5 B.
        In Pilot 5 C, participants could socially learn but only had myopic solutions to choose from.
        In Pilot 5 A, participants learned only individually.
    • Results: Performance in both pilots was mixed. While fewer participants used the loss strategy, some did. This was deemed a failure, as the goal was for no (or very few) participants to discover the loss strategy on their own.
  • Pilot 6 B
    • Setup:
        Redesigned the task and experimental flow to prevent backward induction.
        Made unexplored parts of the networks hidden, reflecting real-world constraints like those in Go.
        Enhanced the social learning phase to highlight the exploration advantage of the machine.
    • Results: Promising results, with 3 out of 10 participants understanding the loss strategy but struggling to execute it.
  • Pilot 7 B
    • Setup: Improved the UI in the social learning phase to:
        Make potential strategies more salient when the teacher shows the solution.
        Emphasize the contrast between participants’ suboptimal solutions and the teacher’s optimal solution.
    • Results: Five out of 10 participants somewhat learned the strategy, but many struggled to reproduce it.
  • Pilot 8 A / B
    • Setup: Redesigned the network task to:
        Remove links from higher levels back to the first level, which tempted participants to avoid losses.
        Ensure losses only occurred when moving up levels, with the largest gains exclusive to the highest level.
    • Results:
        Pilot 8 B: Six out of 10 participants successfully implemented the loss strategy, with three doing so without error.
        Pilot 8 A: Confirmed that participants could not discover the loss strategy on their own.
  • Pilot 9 D
    • Setup: Tested the task with more participants (N = 51) to ensure the strategy was not discoverable.
    • Results: None of the participants discovered the loss strategy.
  • Pilot 9 B
    • Setup: Ran another 10 participants in the B configuration.
    • Results: Only 7 out of 20 participants learned the strategy. Some networks where the loss strategy was not superior to the myopic strategy confused participants.
  • Pilot 10 B
    • Setup: Removed networks where the loss strategy performed worse than the myopic strategy.
    • Results: Nine out of 20 participants imitated the loss strategy without error.
Through this piloting, we calibrated the task such that the optimal strategy’s complexity was not trivial but also not impossible for human minds to discover and feasible to learn. For instance, in an early set of pilot studies, a few participants were able to discover the optimal strategy through backwards induction. To more realistically reflect that backwards induction is prohibitively costly and/or impossible due to the sequential ordering of effects in real-world problems, our final task consisted of networks in which paths were revealed sequentially according to players’ moves. Illustrating the defining influence of learnability, in pilots where the optimal strategy involved incurring a precise number of large losses (e.g., no more or less than three), participants struggled to learn this within the limited time available in our experiment. Our final task thus involved a simplified optimal strategy of taking losses whenever presented with one.
Figure S1: Self-learning page. In the first part of the experiment, participants are on their own to solve the full version of the task. This phase is of different length depending on the generation. In generation 1-4, participant can try themselves on two different networks. The main purpose here is for participants to get an better understanding of the task, before entering the social learning phase. In generation 0, participants have no social learning phase. Correspondingly, for these participants the self-learning phase is extended to a total of 6 networks.
Figure S2: Select demonstrator. Before entering the social learning phase, participants can select one out of five potential demonstrators. The only information guiding participants’ selection is the demonstrators’ average score, which is computed from the points collected during the demonstrators’ demonstration phase. Otherwise, each demonstrator is symbolised identically, with no discrimination between human and machine player. Participants were informed that demonstrators could be humans or machines, however. As an additional comparison to make their choice, participants’ own average performance on the previous two self-learning networks is displayed.
Figure S3: Social learning by observation. Participants see a replay of the demonstrator solving the task. The task is displayed in the same way as during self-play, however the UI is deactivated. For each step of the solution, the next move is indicated as a dashed line.
Figure S4: Social learning by repetition. Participants are exposed to the same network and are asked to repeat the demonstrators’ moves. Participant are receiving +100 points for correct repetitions and -100 points for wrong repetitions. If a wrong move is chosen, the right move is indicated by a dashed line. Participants still have to enact the right move, but do not receive points.
Figure S5: Algorithmic performance during training. We trained three Deep Q-learning algorithms – signified by the three distinct colours – during a time period that was similar to the Individual Learning Phase for human participants (i.e., 315 seconds) for 5,000 episodes on 10,000 unique networks (distinct from the experimental ones). The left y-axis references the solid lines which denote the algorithmic agents’ achieved rewards. The right y-axis and dashed lines denote the achieved maximum level, where the highest level could only be reached by incurring losses to progress. We marked the average values for three heuristic benchmarks on both y-axes: a random agent, a myopic agent always choosing the highest immediate reward, and a loss-seeking agent pursuing losses when possible. During training, the algorithms first discover the myopic strategy, before finally converging on the optimal strategy.
Figure S6: Performance of All 30 Populations: Each population consists of five generations, each with eight players. In human-machine populations, there are three machine players in the first generation. Human players are represented by circles and machine players by squares, with the color of each symbol indicating the player’s average score. In generations following the first, each player selects a single teacher. Arrows depict these teacher-student relationships.
Figure S7: Behavioural Congruency of All 30 Populations: Human players are represented by circles and machine players by squares, with the color of each symbol indicating the fraction of the player’s moves that are congruent with those of the machine. A move is considered congruent if, given the same context (including historical actions), the machine would have made the same move. Arrows depict teacher-student relationships.
Figure S8: Coded Strategies of All 30 Populations: Human players are represented by circles and machine players by squares. A yellow circle signifies that the human player, after social learning, described a written strategy resembling deliberate acceptance of losses. Other strategies are represented by dark violet. Arrows depict teacher-student relationships.
Figure S9: Evolution of Task Performance, Behavioural Congruency, and Strategy Description: Panels A, B, and C display the average task performance, behavioural congruency, and strategy description by human participants within a generation, respectively. Individual populations are shown as dots, with the diamond symbol indicating the grand average per condition. Panels D, E, and F present the top-performing human participants according to the same three metrics. Background dots represent individual populations, and the diamond symbol marks the average of the maximum values per population. Human-machine populations (orange) include a machine player in the first generation, while human-only populations (green) consist solely of humans. Task performance (Panels A and D) refers to the average points earned in demonstration trials. Human-machine behavioural congruency (Panels B and E) measures the fraction of moves where, given the same context (including historical actions), the machine would have made the same move as the human participant. Strategy Description (Panels C and F) assesses whether, after social learning, the human player described a strategy resembling deliberate acceptance of losses (encoded as 1; otherwise, 0). Points are jittered horizontally and vertically to enhance clarity.
Figure S10: Average reward in the final generation. The agent-based simulation replicates our experimental conditions. We systematically vary the discovery rate $d$, representing the probability of discovering the optimal strategy during individual learning, and the social learning rate $l$, which signifies the rate at which the optimal strategy is adopted when observed from a player of the previous generation. The results illustrate the average reward in the final generation of human populations (panels A and C) and human-machine populations (panels B and D). When each agent can socially learn from the best of five potential demonstrators (panels A and B), rarely discovered solutions can be sustained, including those originating from the machine, when social learning rate is relative high. In contrast, such solutions cannot be maintained when the demonstrator is selected randomly (panels C and D).
Figure S11: Alternative social learning structure. We explored alternative social learning structures to assess the sensitivity of our results. In the main analysis, we replicate our experimental condition: each generation consists of eight participants, with three machines in the first generation and five randomly selected demonstrators to choose from. Here, we independently alter each of these parameters. The first panel shows results for a setting with only one machine. In the second panel, we increase the population to 16 agents. In the final panel, agents can choose from all individuals in the previous generation.
Prediction
1a 1b 2a
Condition = Human Population -309.2 -229.1 -404.6
(57.0) (82.4) (59.4)
[418.4-418.4, 200.1-200.1] [383.1-383.1, 55.5-55.5] [524.8-524.8, 286.1-286.1]
Generation -46.5
[93.2-93.2, 2.7-2.7]
Condition:Generation 86.4
[$23.7$, $146.8$]
Intercept 1,843.5 1,814.8 1,886.5
(40.3) (58.3) (42.0)
[$1763.5$, $1918.2$] [$1698.1$, $1925.8$] [$1806.6$, $1975.1$]
Observations 3,840 960 960
Table S1: Model Coefficients Predicting Reward. We present the estimated coefficients along with their standard errors, and 95% confidence intervals.
Prediction
1a 1b 2a
Condition = Human Population -0.884 -0.657 -1.159
(0.138) (0.190) (0.136)
[1.16-1.16, 0.59-0.59] [1.02-1.02, 0.31-0.31] [1.42-1.42, 0.90-0.90]
Generation -0.154
[0.24-0.24, 0.06-0.06]
Condition:Generation 0.210
[$0.09$, $0.33$]
Intercept 1.197 1.036 1.382
(0.098) (0.136) (0.100)
[$1.02$, $1.38$] [$0.80$, $1.31$] [$1.19$, $1.58$]
Observations 38,400 9,600 9,600
Table S2: Model Coefficients Predicting Human-Machine Behavioural Congruency. We present the estimated coefficients along with their standard errors, and 95% confidence intervals.
Prediction
2b
Condition = Human Population -4.180
[6.30-6.30, 3.11-3.11]
Generation -0.503
[0.89-0.89, 0.10-0.10]
Condition:Generation 0.132
[1.40-1.40,$1.41$]
Intercept -0.578
[1.18-1.18, $0.05$]
Observations 960
Table S3: Model Coefficients Predicting Strategy. We present the estimated coefficients along with their standard errors, and 95% confidence intervals.